Your driving a car at a constant speed of 10 m/s. The total mass of the car and its contents its 1200kg. You apply the brakes and come to a smooth stop in 4.0 seconds. While coming to a stop, how large was the force applied by the road on the car?


A. zero N

B. 3000N

C. 4500N

D. 6000N

E. 9000N


Please anwser and explain. Thank you!
Answer #1

change in momentum=F*t

so 1200*10=F*4

so F=3000N

B
Answer #2

Force= change in momentum/time

therefore= 10*1200-0/4sec =3000 N and that will be in the opposite to the direction of motion